# Local stability guarantees for data-driven quadratically nonlinear models

### v1

@@ IV.3 Von Karman vortex street using the enstrophy
IV.4 Lid-cavity flow
In fluid mechanics, the lid-driven cavity is commonly used as a benchmark for testing numerical methods [48]. A viscous fluid is in a confined geometry of a 2D rectangular container, and its motion is driven by the tangential moving of a bounding wall. The fluid motion is governed by the incompressible Navier-Stokes equations. We then apply the model reduction method previously mentioned, resulting in a POD-Galerkin model. In this work, we select the first 6 POD modes to build a data-driven model. Then the extended trapping SINDy algorithm is applied to compare with direct numerical simulation, 6-modes POD-Galerkin model, 16-modes POD-Galerkin model and a naive SINDy model. As shown in Fig. 8, the extended trapping SINDy model outperforms all other presented data-driven models and indicates strong stability compared to all others, even though the $\bm{A}^{S}$ from the extended trapping SINDy is not fully promoted to negative-definite. The $\bm{A}^{S}$ has one positive eigenvalue which is similar as the one in the $50\%$ noisy Lorenz example in Sec. IV.1. Further hyperparameter tuning might facilitate a fully provably stable model but we found empirically this model already works well for the lid cavity flow.
@@ V Conclusion
In Sec. III, based on the trapping SINDy algorithm by Kaptanoglu et al[13], an extended trapping SINDy algorithm was proposed to promote local stability of data-driven quadratically nonlinear models. One of the guarantees on performance needed to use this algorithm is given by the local stability estimates from the theorem proposed in Sec. II.4. The algorithm performs well on tasks such as system identification and simulation with ROMs without requiring the nonlinearities of the dynamical system to be small to have these guarantees. Successfully promoting the local stability of data-driven models requires a combination of well designed algorithms, methods of analysis with physical insight and interpretations, and local or global stability guarantees. The work of this paper was motivated, in part, by the idea that tools that extract models from data should guarantee the stability of models that capture the dominant behavior of bounded physical systems like fluids which can be described by Navier-Stokes equations. We observe many situations where the characteristics of the dynamics or stability/instability conditions are unclear in our presented example results and unpublished experiment. This leads to a number of possible directions for future work that would extend the research presented in this paper.
@@
We train our extended trapping SINDy algorithm on a single trajectory for all of the examples presented in this work. And for most data-driven methods, the performance can be improved by training larger datasets, meaning processing data from multiple trajectories. Modern approaches can effectively address the challenge posed by large datasets, which includes manifold Galerkin projection [49] and autoencoder methods [50, 51, 52], which may also address the Kolmogorov width limitations of linear transformations [53].
@@
As mentioned above in Sec. IV.1 and Sec. IV.4, although the matrix $\bm{A}^{S}$ from Eq. (15b) is not strictly Hurwitz, we observe robust local stability of these SINDy models. The initial conditions that leads to an unstable solution are still not clear in this work. Even though we have given sufficient conditions for attracting trapping regions to exist in quadratically nonlinear systems and conservative estimates about the size of the local stability domain, an unbounded trajectory has not been yet found when the Hurwitz property is slightly broken. The inability to find a $\bm{m}$ such that $\bm{A}^{S}$ is Hurwitz can always be an issue of a suboptimal minima in the optimization problem, but dynamical explanations are required too.
@@
The robust stability despite weakly breaking the Hurwitz property can be dynamically understood because when there is at least one positive eigenvalue the ellipsoid of positive energy growth, $\Omega_{\rho_{-}}$, transitions to a hyperboloid where the $\dot{K}>0$ region stretches to infinity. Schlegel and Noack showed that generically this will be in conflict with global stability for any initial condition, although most trajectories in the $\dot{K}>0$ region may in fact be bounded.
@@
However, this does not imply instability if the energy-preserving nonlinearity constraint is weakly broken, since the system dynamics can easily allow for trajectories starting in the $\dot{K}>0$ region to move away from the origin, but then eventually cross the $\dot{K}=0$ surface and move back towards the origin. Concluding definitively about local stability in this case is very challenging because of the strong dynamical dependence and this difficulty is in fact a motivation for the Hurwitz assumption on $\bm{A}^{S}$ in order to write down a generic stability theorem. Empirically, the stability of these models appears relatively robust to a small positive eigenvalue in $\bm{A}^{S}$; it is only our stability estimates that break down here. Nonetheless, if all eigenvalues of $\bm{A}^{S}$ are positive, and there are energy-preserving nonlinearities, it should be clear that $\dot{K}>0$ is inevitable at distances far enough from the origin, implying instability for trajectories with initial condition in this area.
@@
It roughly follows from this reasoning that if we promote models such that more and more eigenvalues of $\bm{A}^{S}$ become negative, this will contribute to the stability whether or not $\bm{A}^{S}$ is able to become fully Hurwitz. This is precisely our empirical finding in both the local and global versions of the trapping SINDy algorithm.
@@
Future numerical work includes improving the strategy of enforcing a Hurwitz matrix $\bm{A}^{S}$ in the trapping SINDy algorithm [13]. The strategy to enforce a Hurwitz matrix $\bm{A}^{S}$ is to solve $\bm{\xi}$ and $\bm{m}$ separately in each iteration so that each process is convex. However, the Hurwitz loss term is convex composite and inevitably makes the loss function nonconvex. The convergence properties are unclear as well. A clever strategy for enforcing the Hurwitz property was adopted recently in Goyal [15]. Alternatively, with the help of modern powerful computational platforms, the nonconvex problems can be better solved by implementing this method in deep learning. So developing a deep neural network for this method as in Ouala et al. [14] is another promising line of future work.
@@
Lastly, Kramer raises an optimization-based approach to enlarge the estimate of the stability region for a given Lyapunov matrix $\bm{P}$ [16]. From Thm. 2 and Lyapunov function (23a), we can observe that it is possible to control the magnitudes of the eigenvalues of $\tilde{\bm{A}}^{S}$ using different Lyapunov matrices $\bm{P}$. This implies that some $\bm{P}$ can enlarge the the guaranteed stability region by forming $\tilde{\bm{A}}^{S}$ with an even greater largest eigenvalue so that $\rho_{+}$ is larger. Searching for an optimal $\bm{P}$ through data-driven methods as in Goyal et al. [15] and machine learning algorithms is promising to extend the algorithm in this work even further.
@@ Appendix A Sensitivity to perturbations
Now consider a system that is locally or globally bounded by exhibiting a monotonic trapping region as described in the present work. Without nonlinearity, or if the perturbation to the nonlinearity $\delta Q_{ijk}$ is somehow energy-preserving, then the requirement to retain local or global boundedness reduces to the simple requirement that
@@
$\epsilon_{Q}$ $=\|Q_{ijk}+Q_{jik}+Q_{kij}+\delta Q_{ijk}+\delta Q_{jik}+\delta Q_{kij}\|_{\text{max}}$
$=\frac{4\lambda_{1}^{2}(\bm{A}^{S}+\delta\bm{A}^{S})}{3r^{\frac{3}{2}}\|\bm{d}+\delta\bm{d}\|_{2}}.$ (50)
@@
This estimate allows us to roughly categorize quadratically nonlinear dynamical systems into three types: (1) systems very close to a stability threshold $\lambda_{1}\approx 0$ that likely require explicit enforcement of the Hurwitz constraint during optimization because of this sensitivity, (2) systems not close to the $\lambda_{1}\approx 0$ stability threshold, such that e.g. $1+\frac{\delta\lambda_{1}}{\lambda_{1}}\approx 1$, but close to the nonlinear threshold related to $\epsilon_{Q}$, where loss of local stability (estimates) may easily occur, and (3) systems which are robust to general perturbations in $\delta\bm{A}^{S}$ and $\delta\bm{Q}$. It turns out that many low-dimensional systems are relatively robust with regards to the local stability estimates in this work (although the global stability falls apart immediately if $\delta\bm{Q}$ is not enforced to be energy-preserving). For instance, an identified Lorenz model on noisy data has $r=3$, might exhibit $\lambda\approx-1$, $\|\bm{d}+\delta\bm{d}\|_{2}\approx\|\bm{d}\|_{2}=100$ and therefore a modest perturbation,
$\epsilon_{Q}=\|\delta Q_{ijk}+\delta Q_{jik}+\delta Q_{kij}\|_{\text{max}}\approx\frac{4}{300\sqrt{3}}\approx 0.0077,$
@@
Therefore, this example shows that explicit estimates can be made of the sensitivity of a dynamical property (a local trapping region) to errors in nonlinear system identification, and provides evidence that generically the problem of correctly capturing the dynamical stability during data-driven model identification is ill-conditioned for many dynamical systems. Of course, we are always free to impose local stability guarantees, whether or not the underlying data is compatible with the trapping theorems, since this can be seen as a form of model regularization. Then it becomes the role of the optimization to attempt to find a solution that both fits the data well and is compatible with the trapping theorems.

### v3

@@ V Conclusion
In Sec. III, an extended trapping SINDy algorithm was proposed to promote local stability of data-driven quadratically nonlinear models. One of the guarantees on performance needed to use this algorithm is given by the local stability estimates from the theorem proposed in Sec. II.4. The algorithm performs well on tasks such as system identification and simulation with ROMs without requiring the nonlinearities of the dynamical system to be small to have these guarantees. Successfully promoting the local stability of data-driven models requires a combination of well designed algorithms, methods of analysis with physical insight and interpretations, and local or global stability guarantees. The work of this paper was motivated, in part, by the idea that tools that extract models from data should guarantee the stability of models that capture the dominant behavior of bounded physical systems like fluids which can be described by Navier-Stokes equations. We observe many situations where the characteristics of the dynamics or stability/instability conditions are unclear in our presented example results and unpublished experiment, which includes that there exists one positive eigenvalue of $\tilde{\bm{A}}^{S}$ and that no stability guarantee is given by the Thm. 2 in some models but yet to find any unstable solution in neither of these situations. This leads to a number of possible directions for future work that would extend the research presented in this paper.
@@
We train our extended trapping SINDy algorithm on a single trajectory for all of the examples presented in this work. And for most data-driven methods, the performance can be improved by training larger datasets, meaning processing data from multiple trajectories.
@@
As mentioned above in Sec. IV.1 and Sec. IV.4, although the matrix $\tilde{\bm{A}}^{S}$ from Eq. (16b) is not strictly Hurwitz, we observe robust local stability of these SINDy models. The initial conditions that leads to an unstable solution are still not clear in this work. Even though we have given sufficient conditions for attracting trapping regions to exist in quadratically nonlinear systems and conservative estimates about the size of the local stability domain, an unbounded trajectory has not been yet found when the Hurwitz property is slightly broken. The inability to find a $\bm{m}$ such that $\tilde{\bm{A}}^{S}$ is Hurwitz can always be an issue of a suboptimal minima in the optimization problem, but dynamical explanations may be required too.
@@
The robust stability despite weakly breaking the Hurwitz property may be somewhat dynamically understood because when there is at least one positive eigenvalue the ellipsoid of positive energy growth, $\Omega_{\rho_{-}}$, transitions to a hyperboloid where the $\dot{v}>0$ region stretches to infinity. Schlegel and Noack showed that generically this will be in conflict with global stability for any initial condition, although most trajectories in the $\dot{v}>0$ region may remain bounded.
@@
Indeed, generally the system dynamics can allow for trajectories starting in the $\dot{v}>0$ region to move away from the origin, but then eventually cross the $\dot{v}=0$ surface and move back towards the origin. Concluding definitively about local stability in this case is very challenging because of the strong dynamical dependence and this difficulty is in fact a motivation for the Hurwitz assumption on $\tilde{\bm{A}}^{S}$ in order to write down a generic stability theorem. Empirically, the stability of these models appears relatively robust to a small positive eigenvalue in $\tilde{\bm{A}}^{S}$; it is only our stability estimates that break down here. Nonetheless, if some eigenvalues of $\bm{A}^{S}$ are positive, and there are energy-preserving nonlinearities, it should be clear that $\dot{v}>0$ is inevitable at distances far enough from the origin, implying instability for trajectories with initial condition in this area.
@@
It roughly follows from this reasoning that if we promote models such that more and more eigenvalues of $\tilde{\bm{A}}^{S}$ become negative, this will contribute to the stability whether or not $\tilde{\bm{A}}^{S}$ is able to become fully Hurwitz. This is precisely our empirical finding in both the local and global versions of the trapping SINDy algorithm.
@@
Future numerical work includes improving the strategy of enforcing a Hurwitz matrix $\tilde{\bm{A}}^{S}$ during optimization. The current strategy to enforce a Hurwitz matrix $\tilde{\bm{A}}^{S}$ is to solve $\bm{\xi}$ and $\bm{m}$ separately in each iteration so that each process is convex. However, the loss term for making a Hurwitz $\tilde{\bm{A}}^{S}$ inevitably makes the loss function nonconvex. The convergence properties are unclear as well. A clever strategy for enforcing the Hurwitz property was adopted recently in Goyal et al[18]. Alternatively, with the help of modern powerful computational platforms, the nonconvex problems can be better solved by implementing this method in deep learning. So developing a deep neural network for this method as in Ouala et al[17] is another promising line of future work.
@@
Lastly, Kramer raises an optimization-based approach to enlarge the estimate of the stability region for a given Lyapunov matrix $\bm{P}$ [19]. From Thm. 2 and Lyapunov function (24), we can observe that it is possible to control the magnitudes of the eigenvalues of $\tilde{\bm{A}}^{S}$ using different Lyapunov matrices $\bm{P}$. This implies that some $\bm{P}$ can enlarge the the guaranteed stability region by forming $\tilde{\bm{A}}^{S}$ with an even greater largest eigenvalue so that $\rho_{+}$ is larger. Searching for an optimal $\bm{P}$ through data-driven methods as in Goyal et al[18] and machine learning algorithms is promising to extend the algorithm in this work even further.
@@ Appendix A Sensitivity to perturbations
Now consider a system that is locally or globally bounded by exhibiting a monotonic trapping region as described in the present work. Without nonlinearity, or if the perturbation to the nonlinearity $\delta Q_{ijk}$ is somehow energy-preserving, then the requirement to retain local or global boundedness reduces to the simple requirement that (taking $\bm{P}=\bm{I}$ for clarity),
@@
$\epsilon_{Q}$ $=\|Q_{ijk}+Q_{jik}+Q_{kij}+\delta Q_{ijk}+\delta Q_{jik}+\delta Q_{kij}\|_{F}$
$=\frac{3\lambda_{1}^{2}(\bm{A}^{S}+\delta\bm{A}^{S})}{4\|\bm{d}+\delta\bm{d}\|_{2}}.$ (54)
@@
This estimate allows us to roughly categorize quadratically nonlinear dynamical systems into three types: (1) systems very close to a stability threshold $\lambda_{1}\approx 0$ that likely require explicit enforcement of the Hurwitz constraint during optimization because of this sensitivity, (2) systems not close to the $\lambda_{1}\approx 0$ stability threshold, but close to the nonlinear threshold related to $\epsilon_{Q}$, where loss of local stability (estimates) may easily occur, and (3) systems which are robust to general perturbations in $\delta\bm{A}^{S}$ and $\delta\bm{Q}$. It turns out that many low-dimensional systems are relatively sensitive with regards to the local stability estimates in this work. For instance, an identified Lorenz model on noisy data might exhibit $\lambda\approx-1$, $\|\bm{d}+\delta\bm{d}\|_{2}\approx\|\bm{d}\|_{2}=100$ and therefore a modest perturbation,
$\epsilon_{Q}=\|\delta Q_{ijk}+\delta Q_{jik}+\delta Q_{kij}\|_{F}\approx\frac{3}{400}=0.0075,$
@@
Moreover, this example shows that explicit estimates can be made of the sensitivity of a dynamical property (a local trapping region) to errors in nonlinear system identification, and provides evidence that generically the problem of correctly capturing the dynamical stability during data-driven model identification is ill-conditioned for many dynamical systems, i.e. small perturbations in the coefficients make big changes in the dynamical stability considered here. Of course, we are always free to impose local stability guarantees, whether or not the underlying data is compatible with the trapping theorems, since this can be seen as a form of model regularization. Then it becomes the role of the optimization to attempt to find a solution that both fits the data well and is compatible with the trapping theorems.